Y = C + I + G

C = a + b(Y - T)

G = gY


a. Find the equilibrium national income(Y)

b. what further restriction on the parameter is needed for solution to exist









1
Expert's answer
2021-04-29T10:30:37-0400

Solution:

a.). The equilibrium national income (Y) = C + I + G

Y = a + b(Y – T) + Io +  gY


Y = "\frac{I_{o} + a - bT }{1 -b - g}"

 

b.). Further restriction on the parameter needed for the solution to exist.

For the solution to exist, it must be that (b+g) < 1, otherwise, Y; C; and G will be negative, which is impossible. Therefore, 0 < (1 – b – g) < 1